Help me!!!!

Q2. Heat exchanger-NTU method (34 pt). An orange juice is being cooled from 85 °C at a rate of 2000 kg/h in a counter-current heat exchanger by cold water. The juice has a specific heat of 3090 J/(kg.K). The cold water enters the heat exchanger at 8°C at a rate of 2950 kg/h、The overall heat transfer coefficient U-485 w/mK and the area A 4.2 m2. Assuming steady-state conditions, calculate: (a) the heat transfer rate (J/s), (b) the exit juice temperature, and (c) the exit water temperature (and compare to your assumption!).
Answer #1

Ans:

  1. Heat transfer rate = hA(Thot – Tcold)

Here h = Convection coefficient = 485 j/s m2 K

            A = Area = 4.2 m2.

            Thot = 85 0C = 358 K

            Tcold = 8 0C = 281 K

So Rate = 1.568 * 105 j/s

  1. The exit temperature of juice :

Mcpjuice = 0.555 * 3090 = 1714.95 J/s K

NTU = UA/C = 1.1878

Now efficiency = 1 – e-NTU = 0.695

So heat transfer rate = Cjuice(Ti –To) = To = 6.43 0C

  1. The exit temperature of Water

Mcpwater = 0.8194 * 4186 = 3130.0084 j/s K

Heat transfer rate = Cwater(Ti – To) = To = 37.714 0C